OpenStudy (anonymous):

Factor Completely: 36x2 − 25 (6x − 5)(x + 5) (6x + 5)(6x + 5) (6x − 5)(6x − 5) (6x + 5)(6x − 5)

OpenStudy (anonymous):

See you can here use the formula: \[(a^2 - b^2) = (a+b)(a-b)\] Here, a = 6x^2 and b = 5^2 Use this formula you will surely reach your answer...

OpenStudy (anonymous):

is it D?